A wheel is subjectedto uniform angular acceleration about its axisInitially its angular velocity is zero In the first 2 sec it rotates through an angle a1 next 2 sec it rotates through an additional angle a2 the ratio of a2/a1

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi


use the formula

Θ =ωt +1/2 α t2  

now first for 2 sec

a1 = 0 + 1/2 α *4

     =

now angel rotate in next 2 sec = angel rotate in 4 sec  - angel rotate in first 2 sec

                                   a2     = 1/2 *α* 16  - 2α

                                            = 6α

 

so a2/a1 = 3

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin